write a quadratic equation having the given numbers as solutions -4,-1 standered form
Found 2 solutions by stanbon, amoresroy:
Answer by stanbon(75887)

write a quadratic equation having the given numbers
as solutions -4,-1 standard form.
-------
f(x) = (x+4)(x+1) = x^2+5x+4
==================================
Cheers,
Stan H.

Answer by amoresroy(361)

write a quadratic equation having the given numbers as solutions -4,-1 standered form
(x+4)(x+1) = 0
x^2 + 5x + 4 = )
-----------------------
Regards,
Roy